OpenStudy (anonymous):

Please help me to do this problem and give me the steps so that i can do in future by self 9.0-(x/2000)-(38x/19000)*(77/158)=8.6 FIND x

OpenStudy (anonymous):

sir can you give me the ans wer?

OpenStudy (anonymous):

ok then please tell me what to do - what will i do with x/2000

OpenStudy (phi):

9.0-(x/2000)-(38x/19000)*(77/158)=8.6 I would first simplify 38x/19000 by dividing top and bottom by 19 this makes it 2x/1000 Then factor 158 as 2*79 (to make 1000 a 2000) \[ -\frac{x}{2000} - \frac{2x}{2000}\frac{77}{79}= 8.6-9.0 \] multiply the first fraction by 79/79 to get a common denonimator \[ -\frac{x}{2000}\frac{79}{79} - \frac{2x}{2000}\frac{77}{79}= -0.4 \] \[ - \frac{233x}{2000 \cdot 79} = -\frac{2}{5} \] now multiply both sides by -2000*79/233 \[ x= \frac{2}{5}\cdot \frac{2000 \cdot 79}{233} \] \( x= 271 \frac{57}{233} \)

OpenStudy (phi):

Does this make sense?

OpenStudy (anonymous):

sir i got it till -233x/2000*79=-0.4 but in next step you mentioned this -0.4 as 2/5 -why please can you tell me

OpenStudy (phi):

-0.4 is the same as \( -\frac{4}{10} \) or \(- \frac{2}{5} \)

OpenStudy (anonymous):

yes sir - i got it and thanks a lot for the help - is there a simpler way to do it as i keep forgetting the long process - however thanks and a great help

OpenStudy (phi):

There is no really simple way. Maybe if you use a calculator, but I doubt it....